OpenStudy (anonymous):

tan^2(x)-sec^2(-x) simplify the expression

OpenStudy (inkyvoyd):

HINT: start with sin^2 x+cos^2 x=1

OpenStudy (inkyvoyd):

divide both sides by cos^2 x

OpenStudy (anonymous):

x=1?

OpenStudy (inkyvoyd):

No, we don't know x

OpenStudy (anonymous):

im talking about your hint

OpenStudy (anonymous):

no sec²(x)=1+tg²(x)

OpenStudy (anonymous):

oh thats an identity

OpenStudy (inkyvoyd):

rewrite the tangent and secant as sines and cosines

OpenStudy (anonymous):

sec(-x)=sec(x)

OpenStudy (anonymous):

so i get sin^2x/cos^2x - 1/cos^2x?

OpenStudy (inkyvoyd):

\(\Huge \frac{\sin^2x}{\cos^2 x}-\frac{1}{\cos^2x}\)

OpenStudy (inkyvoyd):

yes

OpenStudy (anonymous):

sweet so i was on the right track

OpenStudy (anonymous):

wait factor it out of what?

OpenStudy (inkyvoyd):

\(\frac{1}{\cos^2 x}(\sin^2 x-1)\)

OpenStudy (anonymous):

oh so mult both sides by 1/cos^2

OpenStudy (inkyvoyd):

Well sort of. It's not an equality, so we can't legally do that. We can only factor out the 1/cos^2

OpenStudy (anonymous):

okay thats where i kind of get a bit lost im taking cos out of sin^2x/cos^2x?

OpenStudy (inkyvoyd):

Well I just rewrote the expression. distribute it again, and you can see that it is the same

OpenStudy (anonymous):